OpenStudy (anonymous):

Ok i am having problems here \[\frac{\frac{1}{x+2}-5}{\frac{4}{x}-x}\] i need to simplify

OpenStudy (anonymous):

get common denominator in the numerator and denominator individually, then invert the denominator and multiply and reduce.

OpenStudy (curry):

multiply the top by 4/x - x

OpenStudy (anonymous):

\[\frac{4x-x^2}{x}\]

OpenStudy (anonymous):

is that the bottom?

OpenStudy (curry):

yep

OpenStudy (anonymous):

This is what i'm thinking is simplest \[\frac{\frac{12-4x}{x+2}}{4-x}\]

OpenStudy (amistre64):

\[\frac{\frac{1}{x+2}-5}{\frac{4}{x}-x}=N\] \[\frac{1}{x+2}-5=N(\frac{4}{x}-x)\] \[\frac{1(x+2)}{x+2}-5(x+2)=N(\frac{4}{x}-x)(x+2)\] \[x(1-5x-10)=N(x)(\frac{4}{x}-x)(x+2)\] \[5x^2+9x=N(x^2-4)(x+2)\] \[5x^2+9x=N(x-2)(x+2)(x+2)\] \[5x^2+9x=N(x-2)(x+2)^2\] where are we going with this?

OpenStudy (anonymous):

i'm not sure the dirctions say "Simplify the expression" so i was attempting to simplify

OpenStudy (anonymous):

So why did you set it = to N?